OpenStudy (anonymous):

A car is moving at 20 m/s on a road such that the coefficient of static friction between its tires and the road is 0.80. The minimum distance in which the car can be stopped is? @ryanvarghese12790

OpenStudy (anonymous):

use F = ma Let mass of car = m F(friction) = µN = µmg The resultant force acting on the car is just the friction so ma = -0.80 * m * 9.8 a = -7.84 m/s² [negative because the direction of the acceleration is opposite to the motion] v(f)² = v(i)² + 2as s = [v(f)² - v(i)²] / 2a s = [0 - 20²] / [2 * -7.84] [v(f) = 0 because the car stops] s = 25.5 m